AMSTEX Reference Card

(See the TEX Reference Card for further commands)

Formatting

\pagewidth{⟨dimen⟩}  set page width
\pageheight{⟨dimen⟩}  set page height
\hcorrection{⟨dimen⟩}  move page right
\vcorrection{⟨dimen⟩}  move page down
\flushpar  start a paragraph

with no indentation

\boxed#1  boxed formula
\NoBlackBoxes  omit overfull hbox markers
\comment … \endcomment  unprinted comments
\pageno#1  set page number
\nopagenumbers  turn off page numbering

if not using amsppt style

Fonts

Text Fonts

\rm  roman
\it  italic
\bf  boldface
\sl  slant
\smc  small capitals

Math Mode Fonts

\bold#1  bold letter
\loadbold  load bold math symbols
\boldkey#1  bold keyboard symbol
\boldsymbol#1  bold math symbol (e.g. \alpha)
\Cal#1  caligraphic (script) upper case
\frak#1  German Fraktur
\goth#1  German Fraktur (same as \frak)
\Bbb#1  blackboard bold

\rom#1  Roman

Loading Fonts & Symbols (if not using amsppt style)

\loadmsam  load msam symbol font
\loadmsbm  load msbm symbol font
\UseAMSsymbols  define all symbols from msam, msbm fonts
\newsymbol  define a particular symbol

Changing Font Sizes

\tenpoint  use 10 point fonts
\eightpoint  use 8 point fonts
\dsize  use display size
\tsize  use text size
\ssize  use subscript size
\sssize  use subsubscript size

Macro Definitions

\define\cs{… }  define a control sequence
\redefine\cs{… }  redefine a control sequence
\predefine\newcs{\oldcs}  assign new name to a control seq

\operatorname#1  new operator name
\operatornamewithlimits#1  new operator name with limits

\newsymbol  new symbol from msam, msbm fonts

\begin{matrix}c \\ ◯ \end{matrix} 1998 J.H. Silverman, Permissions on back. v1.3

Send comments and corrections to J.H. Silverman, Math. Dept., Brown Univ., Providence, RI 02912 USA. ⟨jhs@math.brown.edu⟩

Footnotes and Insertions

\footnote#1  footnote

\footnote"*"#1  footnote with specified marker

\topinsert . . . \endinsert  insert at top of page

\midinsert … \endinsert  insert in middle of page

\topcaption#1 . . . \endcaption  caption at top of insert

\botcaption#1 … \endcaption  caption at bottom of insert
\vspace{⟨dimen⟩}  leave vertical space in an insert

Hyphenation

\showhyphens#1  show allowable hyphens
\-  discretionary hyphen
\hyphenation#1  add words to hypenation list

Fractions and Such

\frac#1#2  fraction

\dfrac#1#2  display size fraction
\tfrac#1#2  text size fraction
\fracwithdelims()#1#2  fraction with paren. delimeters
\binom#1#2  binomial coefficient
\dbinom#1#2  display size binomial coefficient
\tbinom#1#2  text size binomial coefficient

\underset#1\to#2  typeset #1 under #2
\overset#1\to#2  typeset #1 over #2

\overbrace#1^#2  overbrace with label above
\underbrace#1_#2  underbrace with label below
\sideset^#1\and^#2\to\bigop  superscripts on side of operator
\cfrac … \endcfrac  continued fraction
\lcfrac … \endcfrac  continued fraction flush left
\rcfrac … \endcfrac  continued fraction flush right

Arrows & Commutative Diagrams

@>#1>#2>  right arrow with labels
@<#1<#2<  left arrow with labels
\CD … \endCD  commutative diagram (don’t use &’s)
@V#1V#2V  down arrow with labels
@A#1A#2A  up arrow with labels
@=  long horizontal = sign
@|  long vertical equal sign
@.  leave out an arrow

\pretend#1\haswidth#2  make arrows longer

Accents

Type  Example  In Math  In Text

hat  â  \hat              \^

expanding hat  _{abc}  \widehat  none

check  ǎ  \check            \v
tilde  ã  \tilde            \~

expanding tilde  _{abc}  \widetilde  none

acute  á  \acute            \'
grave  à  \grave            \`
dot  ȧ  \dot              \D

double dot  ä  \ddot             \"

breve  ă  \breve            \u
bar  ā  \bar              \B

vector  a⃗  \vec  none
cedilla  ç  none  \c

Dimensions

Dimensions are specified as ⟨number⟩⟨unit of measure⟩.

point pt pica pc inch in centimeter cm
m width em x height ex math unit mu millimeter mm
1 pc = 12 pt 1 in = 72.72 pt 2.54 cm = 1 in 18 mu = 1 em

Spacing and Dots

\linebreak  force a line break
\newline  force a new line, old line pushed left
\mathbreak  force line break

\allowmathbreak  allow line break

\-  discretionary hyphen
\.  abbreviation period

\, or \thinspace  thin space

\medspace  medium space

\; or \thickspace  thick space

\! or \negthinspace  negative thin space
\negmedspace  negative medium space
\negthickspace  negative thick space
\quad  quad space
\qquad  double quad space
%  comment line
\  one blank space
\phantom#1  blank space size of #1
\hphantom#1  blank space width of #1, no height
\vphantom#1  blank space height of #1, no width
\smash#1  ignore height and depth
\topsmash#1  ignore height

  \botsmash#1  ignore depth

\mathstrut  strut to help vertical spacing
\smallpagebreak  small space between paragraphs
\medpagebreak  medium space between paragraphs
\bigpagebreak  big space between paragraphs
\pagebreak  force a page break
\nopagebreak  forbid a page break
\newpage  force a page break

fill page with blank space

\hdots  horizontal dots
\vdots  vertical dots
\ddots  diagonal dots
\dots  dots in text or formulas
\ldots  low dots in text or formulas
\cdots  center dots in text or formulas

Miscellaneous Operations

\bmod#1  mod as binary operation
\pmod#1  mod with parentheses
\mod#1  same as pmod, but no parens
\pod#1  parentheses, but no “mod”
\sqrt#1  square root

\root#1\of#2  root

\uproot{⟨number⟩}  move root up/down
\leftroot{⟨number⟩}  move root left/right
\iiint  two integral signs
\iiint  three integral signs
\idotsint  integral signs with dots

AMS Preprint Style

\input amstex
\documentstyle{amsppt}

⟨Preamble Commands⟩

\topmatter

⟨Top Matter Commands⟩

\endtopmatter
\document

⟨Body of Document⟩

\enddocument

Preamble Commands

\TagsOnLeft (default) or \TagsOnRight
\TagsAsText (default) or \TagsAsMath
\NoPageNumbers
\NoRunningHeads
\Monograph
\define

Top Matter Commands

\title . . . \endtitle
\author . . . \endauthor
\affil . . . \endaffil
\address . . . \endaddress
\curraddr . . . \endcurraddr
\email . . . \endemail
\date . . . \enddate
\dedicatory . . . \enddedicatory
\thanks . . . \endthanks
\translator . . . \endtranslator
\keywords . . . \endkeywords
\subjclass . . . \endsubjclass
\abstract . . . \endabstract

\toc … \endtoc (Table of Contents) \leftheadtext#1 (set left headline text) \rightheadtext#1 (set right headline text)

Body of Paper Commands

\specialhead . . . \endspecialhead
\head . . . \endhead
\subhead . . . \endsubhead
\subsubhead . . . \endsubsubhead
\proclaim#1 . . . \endproclaim

\rom#1 (Roman font in proclaim)

\demo#1 . . . \enddemo (proof)

\qed (end of proof marker) \roster … \endroster (roster of listed items)

\item (start a new item in a roster) \item[⟨number⟩] (specify roster item number) \item"*" (item with specified marker) \therosteritem#1 (refer to specified roster item) \widestnumber\item#1 (set width for roster labels)

\nofrills (turn off automatic font, spacing, punctuation) \usualspace (usual space following punctuation)

\definition#1 . . . \enddefinition
\example#1 . . . \endexample
\remark#1 . . . \endremark

\block … \endblock (indented text) \cite (cite a reference)

AMS Preprint Style — References

\Refs … \endRefs  list of references
\refstyle#1  specify style A, B, or C

A = initials, B = name, C = number

\ref … \endref  individual reference
\no or \key  number or key for reference

  \widestnumber\no#1 or \widestnumber\key#1

\by  author
\bysame  same as previous author
\paper  name of paper
\vol  volume
\yr  year of publication
\jour  journal

\page or \pages  page(s)
\toappear  to appear

\inbook  article in a book
\moreref  additional reference information
\paperinfo  extra information after paper title
\procinfo  information about proceedings
\issue  issue number
\lang  language
\transl  information about translated version

\book  book

\ed or \eds  editor(s)
\publ  publisher
\publaddr  publisher address
\bookinfo  extra information after book title
\finalinfo  extra information for end

\miscnote  same as \finalinfo, in parens.

Overlines and Underlines

\underline#1  underline
\overline#1  overline

\overarrow#1  over right arrow
\underarrow#1  under right arrow

\overleftarrow#1  over left arrow
\underleftarrow#1  under left arrow
\overleftrightarrow#1  over left-right arrow

Delimeters

[ \lbrack or \[ { \lbrace or \{ ⟨ \langle
] \rbrack or \] } \rbrace or \} ⟩ \rangle
| \vert or \|  ⌊ \lfloor  ⌈ \lceil
∥ \Vert or \|  ⌋ \rfloor  ⌉ \rceil
↑ \uparrow  ⇑ \Uparrow  ↕ \updownarrow
↓ \downarrow  ⇓ \Downarrow  ⇕ \Updownarrow
[[ [\![  (( (\!(  ⟨⟨ \langle\!\langle
]] ]\!]  )) )\!)  ⟩⟩ \rangle\!\rangle

\left#1 \right#1  expanding delimeters
\left. \right.  empty delimeters

\bigl#1 \bigr#1  big delimeters

\Bigl#1 \Bigr#1  bigger delimeters
\biggl#1 \biggr#1  even bigger delimeters

Non-Italic Function Names

\arccos  \cos   \csc  \exp  \ker     \limsup  \min  \sinh
\arcsin  \cosh  \deg  \gcd  \lg      \ln      \Pr   \sup
\arctan  \cot   \det  \hom  \lim     \log     \sec  \tan
\arg     \coth  \dim  \inf  \liminf  \max     \sin  \tanh

Alignments and Displayed Equations

\\  separate lines
&  separate items in a line
\align … \endalign  align equations, full

width of page

                                                               \alignat#1 . . . \endalignat  align #1 pairs
                                                                 \xalignat#1 . . . \endxalignat  equally spaced
                                                                 \xxalignat#1 . . . \endxxalignat  equally spaced, flush

\aligned … \endaligned  align equations, width

as needed

                                                                 \alignedat#1 . . . \endalignedat  align #1 pairs
                                                                 \topaligned . . . \endtopaligned  align along top
                                                                 \botaligned . . . \endbotaligned  align along bottom

\gather … \endgather  centered equations, full

width of page

                                                               \gathered . . . \endgathered  centered equations,

width as needed

\multline … \endmultline  first line left, middle lines

centered, last line right

\shoveleft#1  shove lines left
\shoveright#1  shove lines right
\multlinegap{⟨dimen⟩}  change margins
\cases … \endcases  case construction
\split … \endsplit  align split equations with

variable tag placement

\Sb … \endSb  multi-line subscript
\Sp … \endSp  multi-line superscript
\text#1  text within formula
\intertext#1  text between lines
\foldedtext#1  lines of text in formula
\topfoldedtext#1  top-aligned folded text
\botfoldedtext#1  bottom-aligned folded text
\foldedwidth{⟨dimen⟩}  set width of folded text
\allowdisplaybreak  allow page break after line
\allowdisplaybreaks  allow page breaks after any line
\displaybreak  force page break after line
\vspace{⟨dimen⟩}  extra space between two lines
\spreadlines{⟨dimen⟩}  extra space between every line

\spreadmatrixlines{⟨dimen⟩} same for a matrix

\jot  unit of vertical space
\tag#1  tag for a formula
\thetag#1  refer to tag in current style
\tag"*"  tag exactly as specified

Matrices

\matrix … \endmatrix  matrix alignment
\pmatrix … \endpmatrix  matrix with parentheses
\bmatrix … \endbmatrix  matrix with brackets
\vmatrix … \endvmatrix  matrix with vertical lines
\Vmatrix … \endVmatrix  matrix with double vertical lines

                                                               \smallmatrix . . . \endsmallmatrix small matrix

\format  specify a format for a matrix
\c \l \r  format entry center, left, right

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1998 J.H. Silverman, November 1998 v1.3 Math. Dept., Brown Univ., Providence, RI 02912 USA TEX and AMSTEX are trademarks of the American Mathematical Society Permission is granted to make and distribute copies of this card provided the copyright notice and this permission notice are preserved on all copies.